# A Topological Gaussian Mixture Model for Bone Marrow Morphology in Leukaemia Qiquan Wang1,†, Anna Song1,2, Antoniana Batsivari2, Dominique Bonnet2, and Anthea Monod1

### v1

@@ 5.2 Model Evaluation
PH1 NW.
We begin our evaluation by studying the NW quadrant which characterizes the number of loops, their thickness, and sizes, and was seen to give rise to distinguishing features in the samples’ estimated density between the different phases. On NW samples, taking into account the trade-off between the model fit, computational time, and ensuring that the model is generic across all the samples within one phase, we found that the optimal number of components for the Phase O model to be 3, then increasing the number of components for each consecutive phase by 1 was found to distinguish the models sufficiently as well as allow good fits to the samples within each phase. Figure 12 shows the GMMs for each phase and interpreting from the parameter of the GMMs, we found that compared to the CTRL samples, the Phase I samples tend to contain thinner and more compact vessels in the early stages of AML engraftment. Transitioning from Phase I to II, towards later stages of engraftment, we observe instead a shift in density towards the northwest direction in the diagram, in favor of higher values of and sizes, indicative of thicker and wider loops formed by vessels in the cavity, although not as much as those found in Phase O.
We obtain our parameter estimates for each phase model over 50 bootstrap iterations on the training set, and we evaluate these models on 50 bootstrap samples from the test set. For evaluation. we measure the similarities between the phase model and the sample density approximates using the Hellinger distance (Hellinger,, 1909), computed as
between two densities and . Figures 13(a), 13(b), and 13(c) show boxplots of the Hellinger distance computed between density approximations of bootstrap samples from each of the NW samples with each of the phase models respectively. We see from Figure 13(a) that the approximate density for the control samples, colored in green, are on average most similar to the Phase O model. Similarly, in Figure 13(b), we find that on average the densities of the Phase I samples, denoted in orange, are on average most similar to the Phase I model, except for the HL60 at 25%, P1 at 76%, and MNC at 53% samples. Lastly, Figure 13(c) shows that the densities of the Phase II samples are on average most similar to the Phase II model as indicated by the box and whisker plots in red. A further quantitative evaluation by taking the sum of the distances can be found in Appendix E.2 which supports the previous claims.
PH1 NE.
We next study the neighboring quadrant PH1 NE. From the transitions between the Phase O, I, and II models for NE shown in Figure 14, we find that compared to the unaffected bone marrow vasculature in the CTRL samples, samples in the early stages of AML engraftment have more compact vasculature. We observe a higher probability around lower values of both and critical sizes, suggesting tighter spacing between vessels in general, both in the narrowing between loops and loops formed by vessels. Then, transitioning from the Phase I model to the Phase II model, we observe the region of high probability moving towards higher and values, suggesting a dilation of the underlying vessel structure. Again, Figure 30 of the relative Hellinger distances between the models and approximations shows that on average most of the Phase I and II samples agree with their respective phase models.
PH2 NE.
Evaluating our model on the NE quadrant. Between the Phase O and I models, we again find a shift in the region of high probability towards lower values of and , indicating smaller loops and cavities formed by vessels. This suggests a more compact vasculature at the onset of angiogenesis. As we transition from the Phase I to the Phase II model, we find that the distribution shifts towards larger values of and , indicating that the vasculature becomes more sparse. Figure 15 shows the composition of the Gaussian mixture model and densities plotted as heatmaps. Figures for the evaluation of the model can be found in Appendix E.
The collection of points in the NW quadrant can be disregarded since they arise as a complication in the piecewise computation of SDPH along the bone images. In theory, there should only be one point with the quadrant for each sample representing the entire connected vessel structure. The points in the NW can also be disregarded since it contains only a few points which should never, or only very rarely, occur since they correspond to bubbles of extravascular space trapped inside the lumen and may only appear due to noise. The evaluation for the SW and SW quadrant can be found in the Appendix E as the type of characteristics described within these quadrants are not distinct between the phases.
Prediction.
We can also use the estimated phase models to predict the most likely phase of a “new” sample. We do this by finding the most similar phase model to its GMM approximation measured by the Hellinger distance. To evaluate the predictive accuracy of this method of classification, we take bootstrap samples of the test set from each bone sample and independently derive their labels. Over 50 bootstrap samples of each of the test datasets, we obtain accuracies of 89.5 on the NW bootstrap samples and 76.1 on the NE bootstrap samples. Figure 16 shows breakdowns of the accuracies for each bone sample based on the NW quadrant and the NE quadrant. Note that the inaccurate prediction for the HL60 sample at 23% engraftment and the P1 sample at 40% is to be expected, since in both the global and local texture analyses they were also misclassified and seen as anomalous. Given the level of engraftment relative to the rest of the samples within Phase I, the MNC sample at 53% engraftment also appears to be anomalous.
@@ 6 Discussion
In this paper, we presented a new topological method for classifying and modeling the architectural changes in the bone marrow vasculature associated with the progression of AML. We quantified these changes using persistent homology applied to confocal microscopy images. Specifically, we implemented SDPH to quantify and analyze the data, enabling us to categorize data samples into morphologically-derived phases corresponding to different stages of disease progression. Our analysis reveals that clustering based on SDPH summaries aligns closely with biological intuition regarding the stages of AML progression, indicating that SDPH clustering is an effective method for studying AML. We subsequently developed a class of phase-dependent GMMs that effectively approximated the spatial patterns captured within the SDPH diagrams, reflecting transitions in vascular morphology. We have shown that these models can be employed to classify new observations into phases or stages of AML progression.
@@
Our findings indicated that the NW quadrant of the SDPH diagrams for the knee regions is notably more informative and discriminative across stages of engraftment than other regions. This is reflected in both the accuracy of our models and their predictive performance. This observation aligns with the understanding that points in this quadrant correspond to “real” textures, representing static cycles in the vascular network, rather than “virtual” textures (noise arising from the construction of the filtration of the signed distance function). Thus, modeling on the NW region alone provides sufficient information about the vessel structure to allow for accurate prediction of the disease progression.
Biological Implications.
@@
While most previous studies have primarily focused on investigations of late-stage leukaemia morphology (Passaro et al.,, 2017; Duarte et al.,, 2018), our work successfully distinguishes between the early and late stages of AML progression, highlighting clear morphological divergence between them quantified using SDPH. The morphological features identified by our models and analyses correspond to known biological processes: in the early stage, we identified higher number of thinner, more compact vessels as a result of angiogenesis; in the late stage, vessels are sparser apart due to clusters of AML cells deforming the bone marrow causing spatial constraints. Given these distinct differences between early and late stages, our work provides a solid, quantitative foundation for further research into the biological mechanisms driving these patterns, particularly during the early stages of AML progression, and the development to later stages.
Limitations and Future Outlook.
@@
Our study was subject to notable constraints, given the nature of the problem of interest and the data available. These are a limited dataset, which necessitated the use of subsampling methods for classification; and the inherent uncertainty due to the levels of engraftment having been derived as global averages of bones other than the imaged femur, which arises in the reported percentage, and thus prevents their use as exogenous variables.
@@
There are important and interesting future directions of research in the statistical and computational components of our study, in addition to the previously-discussed biological component. Given a larger collection with more continuous measures of engraftment levels or time after engraftment, we envision extending our model to a “time”-dependent Gaussian mixture model for the progression, where the parameters of the GMM vary with time or engraftment level, rather than fixed for each phase. This enhancement could lead to more accurate predictions of AML progression. We also envision that our framework could be adapted for application to other biological systems with similar vascular structures.
@@
The raw imaging data from this study is proprietary data and cannot be openly shared, however, the summaries of the analysis such as the vectorized summaries, SDPH diagrams, and code to replicate steps of the methodology however can be found at this repository.
We utilized the following packages in Python for the computation of SDPH: giotto-tda (Tauzin et al.,, 2020); GUDHI (The GUDHI Project,, 2015); and CubicalRipser (Kaji et al.,, 2020). Clustering was performed using the scikit-learn package in Python (Pedregosa et al.,, 2011). The KeOps package (Charlier et al.,, 2021) was used for the efficient computation of Gaussian kernel estimates. The pomegranate package (Schreiber,, 2018) was used for the training and evaluation of the models. A comprehensive list of packages used can be found in the above repository.
@@ Acknowledgments
Q.W. is funded by a CRUK–Imperial College London Convergence Science PhD studentship at the UK EPSRC Centre for Doctoral Training in Modern Statistics and Machine Learning (2021 cohort, PIs Monod/Williams), which is supported by Cancer Research UK under grant reference [CANTAC721\10021]. A.S. was funded by the Francis Crick Institute and Imperial College London. A.B. was funded by an junior EHA fellowship. A.M. is supported by the Engineering and Physical Sciences Research Council under grant reference [EP/Y028872/1].
@@ Appendix A Tri-Cluster Texture Decomposition of the Long Regions of Bone Samples
In Section 4.2, we presented the texture decomposition of the knee region of the bone sample with three clusters. It was noted that applying the same technique on the long regions of the bone samples would not yield similar results due to the structural complexity along the length of the femur bone. Here we present the exact comparison on the long region of the bone when we apply the same texture decomposition with three clusters. Figure 17 shows the three clusters projected onto 2D slices of the long bone samples. We label the white, light blue, and dark blue colored textures as textures A, B, and C. We then see that, similar to the knee regions of the sample, the CTRL samples are mainly composed of texture A, however, this behavior is also seen in other samples such as the P1 samples and U937 sample at 1% engraftment. The introduction of a large proportion of the darker texture C distinguishes the samples with low levels of engraftment from the CTRL samples. Transitioning to the samples with higher levels of engraftment, we observe diminishing proportions of texture C replaced by largely texture B and texture A. Yet, we see visible similarities between the CTRL samples and samples with medium levels of engraftment (5̃5%), which cannot be explained by the patterns identified in the knee region and may be due to the highly nonlinear spread of the AML cells in the bone.
Figure 18(a) shows the proportions of each texture present in each sample, where we apply PCA to the decomposition, embedding the texture decomposition of each sample into the plane spanned by the first 2 principal components of the reduction. Figure 18(b) displays this embedding. Here, the boundaries are less discernible; in particular, the second CTRL sample presents itself as more similar to the samples with high levels of engraftment and the P1 60% sample is found to be “closer” to the CTRL samples and samples with much lower levels of engraftment, despite some similarities to those behaviors identified in the knee region. Furthermore, this finding is consistent with higher numbers of clusters experimented.

### v2

@@ 5.3 Model Evaluation
We evaluate the classification performance of our phase models using leave-one-out cross-validation (LOOCV) across all 27 samples under two regimes. In the first, standard LOOCV is performed directly on the original samples. Due to data imbalance, particularly the sparsity of Phase O samples, this setup can lead to biased or unstable predictions. In the second regime, we augment the training set using bootstrap resampling: within each LOOCV iteration, we concatenate training samples within each phase and generate 50 bootstrap subsamples per phase. This augmentation ensures balance across phases samples and allows for more stable estimation of the phase-dependent GMM parameters.
For each test sample, we compute its similarity to the fitted Phase O, I, and II models using two measures: the Hellinger distance and the symmetrized KullbackāLeibler (KL) divergence. The test sample is then assigned to the phase corresponding to the model with the smallest distance or divergence to its density GMM approximation. The Hellinger distance (Hellinger,, 1909) between densities and is defined as
while the KullbackāLeibler (KL) divergence (Kullback & Leibler,, 1951) is given by
As the KL divergence is asymmetric, we report the symmetrized version by averaging and .
Method | PH1_NE | PH0_SW | PH1_NW | PH1_SW | PH2_NE |
Hellinger-based classification | |||||
LOOCV | 0.704 | 0.482 | 0.852 | 0.370 | 0.889 |
LOOCV & bootstrap | 0.852 | 0.556 | 0.852 | 0.296 | 0.889 |
KL-divergence-based classification | |||||
LOOCV | 0.778 | 0.482 | 0.852 | 0.185 | 0.889 |
LOOCV & bootstrap | 0.852 | 0.556 | 0.852 | 0.269 | 0.889 |
TableĀ 4 summarizes the classification accuracies under both LOOCV regimes and distance metrics, evaluated over 5 representative quadrants. We observe high performance in the PH1 NE, PH1 NW, and PH2 NE quadrants, with consistent improvements under bootstrap. In particular, classification accuracies under LOOCV with bootstrap reach 85.2%, 85.2%, and 88.9% for these 3 regions respectively, with PH2 NE yielding the highest overall accuracy. These results suggest that quadrant-specific models effectively capture structural differences in bone marrow morphology across the main stages of AML engraftment and are sufficient for reliable phase discrimination.
In contrast, classification performance is lower in the PH0 SW and PH1 SW regions. This aligns with earlier model interpretations where it was noted that these quadrants displayed greater structural ambiguity and less distinct separation between phases. A recurring source of misclassification is the P1 sample at 40% engraftment, which was also flagged in the local and global analysis as a potential boundary case. This reflects the biological variability and overlap between phases, particularly at transitional stages of engraftment.
5.4 Method Comparison
We now compare the performance of our phase-dependent GMM applied to SDPH diagrams with several alternative approaches for phase classification, in order to assess the added value of our proposed framework over existing topological and morphological methods. Specifically, we consider: (i) a phase-dependent GMM applied to distance persistent homology (DPH) diagrams; (ii) classification algorithms such as -nearest neighbors (-NN) and support vector machines (SVM/SVC) trained on SDPH landscapes (Bubenik,, 2015); (iii) classification on persistence images (Adams etĀ al.,, 2015) computed from SDPH diagrams; and (iv) a baseline based on traditional morphometric features.
Gaussian Mixture Models on Distance Persistent Homology. We compute persistent homology based on distance function (1) from Section 3.2, which tracks the distance of a point to the surface boundary of a set , without any sign indicating whether the point lies inside or outside the set as in (2); we refer to this as distance persistent homology (DPH). The resulting DPH persistence diagrams consist of the absolute values of the birth and death points in the SDPH diagrams. Note that this approach corresponds to computing classical distance-based persistent homology, which is the typical strategy in most application areas seeking to apply TDA tools.
To fit the phase-dependent GMM onto the DPH diagrams, we optimized the mixture by balancing computational efficiency with the BIC, also ensuring that no Gaussian component had zero weight. The optimal mixture sizes for the and diagrams for Phases O, I, and II were 3, 4, and 5, respectively. For the diagrams, the optimal mixture sizes for Phases O, I, and II were 4, 2, and 2, respectively. For a quantitative comparison of model performance, we evaluated the classification performance of these models using LOOCV with phase-level bootstrap augmentation. Under this regime, the predictive accuracies achieved using the Hellinger distance were 66.7% for PH0, 70.4% for PH1, and 81.5% for PH2 diagrams. The corresponding results using the KL-divergence were 63.0%, 70.4%, and 85.2%. The PH0 and PH1 results are considerably lower than those achieved by the corresponding SDPH-GMM models, reinforcing the added value of incorporating signed distance information. While standard DPH retains the topological summary of the data, it fails to capture the nuanced directional and morphological differences that SDPH encodes. These findings highlight the advantage of our model over a direct application of standard persistent homology. Note, the high performance of the PH2 DPHāGMM models is to be expected. In a prior analysis, the PH2 NE quadrant exhibited strong separation between phases, and the full PH2 diagram differs from it only by the inclusion of the PH2 NW region, which introduces mild noise. Nonetheless, our SDPHāGMM still achieves highest accuracy (88.9%), demonstrating the overall superiority of our framework.
Classification on SDPH Landscapes. The algebraic construction of persistence diagrams makes them difficult to integrate into standard statistical and machine learning methodology, due to the complicated geometry that they induce (Turner etĀ al.,, 2012). A popular approach to bypassing this problem is to apply a vectorization method; one of the most popular vectorizations is the persistence landscape (Bubenik,, 2015). Briefly, landscapes consider the convex hull of persistence points over the diagonal rotated to the -axis, giving a functional summary of the persistence diagram which is feasible to work with in statistics and machine learning. This comparison experiment studies the performance of another standard approach to using persistent homology.
We evaluate this again under LOOCV with bootstrap. We converted the resultant PH0, PH1, and PH2 diagrams as well as the PH1 NW, PH1 NE, and PH2 NE quadrants in each subsample into persistence landscapes. After training and testing using 4 classifiersālogistic regression, SVM, random forest, and -NNāwe observed varied accuracies depending on the specific diagram subset and classifier as shown in Table 5. The highest accuracy of 63.0% was achieved by random forest when using all dimensions of the landscapes, while accuracies for other combinations ranged from as low as 22.2% (logistic regression on PH1 diagrams and -NN using all dimensions) to around 59.3% (random forest on PH2 NE quadrant). These results indicate that our proposed signed distance method clearly outperforms this standard vectorization approach and that the persistence landscape construction procedure applied to SDPH diagrams results in a significant loss of important information for accurate inference.
Classification on SDPH Images. As an alternative vectorization scheme to persistence landscapes, persistence images preserve the importance of feature persistence while enabling the use of standard machine learning algorithms (Adams etĀ al.,, 2015). Each point in the persistence diagram is mapped to the birth-persistence plane where it contributes a localized Gaussian kernel typically weighted by persistence allowing more persistent features to exert greater influence. The resulting continuous function is discretized over a grid to form a pixel-based image.
Similar to the landscapes, we compute the persistence images for the PH0, PH1, and PH2 diagrams as well as the PH1 NW, PH1 NE, and PH2 NE quadrants. Under the same LOOCV and bootstrap regime, with logistic regression, SVM, random forest, and -NN classifiers, persistence images consistently outperformed landscapes. As shown in Table 5, the highest accuracy achieved was 81.5% using logistic regression on PH1 (NE) diagrams. Overall, accuracies ranged from 48.1% to above 70%, indicating improved discriminatory power relative to persistence landscapes. While this performance is competitive with our proposed model, which achieved a highest accuracy of 88.9%, our goal is not only to attain high classification accuracy but also build an interpretable, parameterized model that captures and explains the progression of AML. This is achieved through our phase-dependent GMM framework grounded in the theoretical results of Song etĀ al., (2025), which lend interpretability to the possible shapes and sizes of vessels and gaps within the bone marrow vasculature over the course of AML progression.
Classifier | All Dims | PH0 | PH1 | PH2 | PH1 NW | PH1 NE | PH2 NE |
Persistence Landscapes | |||||||
Logistic Regression | 0.556 | 0.556 | 0.222 | 0.519 | 0.296 | 0.333 | 0.481 |
SVM | 0.370 | 0.444 | 0.407 | 0.556 | 0.296 | 0.296 | 0.481 |
Random Forest | 0.630 | 0.407 | 0.222 | 0.556 | 0.222 | 0.444 | 0.593 |
-NN | 0.222 | 0.481 | 0.259 | 0.519 | 0.296 | 0.296 | 0.481 |
Persistence Images | |||||||
Logistic Regression | 0.704 | 0.630 | 0.815 | 0.741 | 0.815 | 0.778 | 0.741 |
SVM | 0.667 | 0.556 | 0.630 | 0.630 | 0.667 | 0.630 | 0.630 |
Random Forest | 0.741 | 0.481 | 0.704 | 0.741 | 0.630 | 0.741 | 0.704 |
-NN | 0.667 | 0.556 | 0.630 | 0.630 | 0.667 | 0.630 | 0.630 |
Classification on Morphometrics. To contextualize the difficulty of phase classification using morphometric measures, we present a classification task on a simplified dataset. Specifically, we selected a representative sample from each of the Phase O, I, and II groups of samples and extracted 30 subregions from each for classification. For each subregion, we computed a comprehensive set of morphometric features derived from the literature (Brown etĀ al.,, 2022; Lafage-Proust etĀ al.,, 2015; Kvasnicka & Thiele,, 2004), encompassing vessel quantity, quality, network architecture, and spatial relationships. These included measures such as total vessel volume, vessel number per unit area, mean vessel diameter, curvature, tortuosity, total vessel length, junction counts, and inter-vessel distances.
We then trained 3 standard classifiersārandom forest, logistic regression, and SVMāon these features to discriminate between the 3 phases. The resulting classification accuracies were 66.7% (random forest), 64.4% (logistic regression), and 54.4% (SVM), indicating limited discriminative power. These results highlight the intrinsic difficulty of the task, especially when relying solely on conventional morphometric descriptors. Singular morphometric measures often fail to capture the complex, large-scale heterogeneity of bone marrow architecture, as evidenced by the poor classification performance in this setting. The overlapping and variable nature of vascular structures across phases likely contributes to the limited separability, underscoring the need for more expressive and topology-aware representations such as those employed in our framework.
@@ 6 Discussion
In this paper, we presented a new topological method for classifying and modeling the architectural changes in the bone marrow vasculature associated with the progression of AML. We quantified these changes using persistent homology applied to proprietary confocal microscopy images where the fineness of the resolution is a novelty in AML imaging. Specifically, we implemented SDPH to quantify and analyze the data, enabling us to categorize data samples into morphologically-derived phases corresponding to different stages of disease progression. Our analysis reveals that clustering based on SDPH summaries aligns closely with biological intuition regarding the stages of AML progression, indicating that SDPH clustering applied to images is is an effective method to study AML. We subsequently developed a class of phase-dependent GMMs that effectively approximated the spatial patterns captured within the SDPH diagrams, reflecting transitions in vascular morphology. We have shown that these models can be employed to classify new observations into phases or stages of AML progression. Through a comparison with straightforward approaches that compute classical persistent homology or apply a standard vectorization, we show that on fine resolution confocal microscopy images, the SDPH framework is required for more meaningful inference results.
@@
Our findings indicated that the NE and PH1 NW quadrants of the SDPH diagrams are notably more informative and discriminative across stages of engraftment compared to other regions. This is reflected in both the accuracy of our models and their predictive performance. This observation aligns with the understanding that points in this quadrant correspond to ārealā textures, representing static cycles in the vascular network, rather than āvirtualā textures (noise arising from the construction of the filtration of the signed distance function). Thus, modeling on the NE or PH1 NW regions alone provides sufficient information about the vessel structure to allow for accurate prediction of the disease progression.
Biological Implications. While most previous studies have primarily focused on investigations of late-stage leukaemia morphology (Passaro etĀ al.,, 2017; Duarte etĀ al.,, 2018), our work successfully distinguishes between the early and late stages of AML progression, highlighting clear morphological divergence between them quantified using SDPH. In addition to our SDPH approach combined with GMMs for inference, an additional biological novelty of our work is the study of early vs.Ā late stage in a comparative context, while previous significant AML research interests have focused on studying primarily the late stage. Our study was possible due to the quality of our proprietary confocal microscopy imaging dataset at such fine resolutions.
@@
The morphological features identified by our models and analyses correspond to known biological processes: in the early stage, we identified higher number of thinner, more compact vessels as a result of angiogenesis; in the late stage, vessels are sparser apart due to clusters of AML cells deforming the bone marrow causing spatial constraints. Given these distinct differences between early and late stages, our work provides a solid, quantitative foundation for further research into the biological mechanisms driving these patterns, particularly during the early stages of AML progression, and the development to later stages.
@@
Limitations and Future Outlook. Our study was subject to notable constraints, given the nature of the problem of interest and the data available. Our data are limited, given the high time and equipment costs required to generate them, which necessitated the use of subsampling methods for classification. Another limitation is the inherent uncertainty due to the levels of engraftment having been derived as global averages of bones other than the imaged femur, which arises in the reported percentage, and thus prevents their use as exogenous variables.
@@
There are important and interesting future directions of research in the statistical and computational components of our study, in addition to the previously-discussed biological component. Given a larger collection with more continuous measures of engraftment levels or time after engraftment, we envision extending our model to a ātimeā-dependent Gaussian mixture model for the progression, where the parameters of the GMM vary with time or engraftment level, rather than fixed for each phase. This enhancement could lead to more accurate predictions of AML progression. We also envision that our framework could be adapted for application to other biological systems with similar vascular structures. Moreover, our model naturally lends itself to extensions for uncertainty quantification, for example, by constructing confidence regions around GMM components to assess compatibility with a given phase model. This, in turn, may help with identifying atypical or borderline sample behaviors arising from biological variationāpatterns we have already begun to observe in our analysis.
@@
The raw imaging data from this study is proprietary data and cannot be openly shared, however, the summaries of the analysis such as the vectorized summaries, SDPH diagrams, and code to replicate steps of the methodology can be found at this repository.
We utilized the following packages in Python for the computation of SDPH: giotto-tda (Tauzin etĀ al.,, 2020); GUDHI (The GUDHI Project,, 2015); and CubicalRipser (Kaji etĀ al.,, 2020). Clustering was performed using the scikit-learn package in Python (Pedregosa etĀ al.,, 2011). The KeOps package (Charlier etĀ al.,, 2021) was used for the efficient computation of Gaussian kernel estimates. The pomegranate package (Schreiber,, 2018) was used for the training and evaluation of the models. A comprehensive list of packages used can be found in the above repository.
@@ Acknowledgments
Q.W.Ā is funded by a CRUKāImperial College London Convergence Science PhD studentship at the UK EPSRC Centre for Doctoral Training in Modern Statistics and Machine Learning (2021 cohort, PIs Monod/Williams), which is supported by Cancer Research UK under grant reference [CANTAC721\10021]. A.S.Ā was funded by the Francis Crick Institute and Imperial College London. A.B.Ā was funded by an junior EHA fellowship. A.M.Ā is supported by the Engineering and Physical Sciences Research Council under grant reference [EP/Y028872/1].
@@ Appendix A Tri-Cluster Texture Decomposition of the Long Regions of Bone Samples
In Section 4.2 of the main paper, we presented the texture decomposition of the knee region of the bone sample with 3 clusters. We note that applying the same technique to the long region of the femur bones may not yield comparable results, given the greater structural variability and complexity observed along the boneās length. Here we present the exact comparison on the long region of the bone when we apply the same texture decomposition with 3 clusters. Figure 16 shows the 3 clusters projected onto 2D slices of the long bone samples. We label the white, light blue, and dark blue colored textures as textures A, B, and C. We then see that, similar to the knee regions of the sample, the CTRL samples are mainly composed of texture A, however, this behavior is also seen in other samples such as the P1 samples and U937 sample at 1% engraftment. The introduction of a large proportion of the darker texture C distinguishes the samples with low levels of engraftment from the CTRL samples. Transitioning to the samples with higher levels of engraftment, we observe diminishing proportions of texture C replaced by largely texture B and texture A. Yet, we see visible similarities between the CTRL samples and samples with medium levels of engraftment (55%), which cannot be explained by the patterns identified in the knee region and may be due to the highly nonlinear spread of the AML cells in the bone.
Figure 17(a) shows the proportions of each texture present in each sample, where we apply PCA to the decomposition, embedding the texture decomposition of each sample into the plane spanned by the first 2 principal components of the reduction. Figure 17(b) displays this embedding. Here, the boundaries are less discernible; in particular, the second CTRL sample presents itself as more similar to the samples with high levels of engraftment and the P1 60% sample is found to be ācloserā to the CTRL samples and samples with much lower levels of engraftment, despite some similarities to those behaviors identified in the knee region. Furthermore, this finding is consistent with higher numbers of clusters experimented.